9, IF the engine of a car of mass m is capable of delivering power P, the minimum time needed for the car to accelerate from rest to a speed v is?


1
Expert's answer
2021-09-05T17:51:01-0400

Gives

Power =p

Mass=m

Speed=v

Energy change

"∆E=\frac{1}{2}mv^2-0"

"∆E=\frac{1}{2}mv^2"

"Power (p)=\frac{W}{t}=\frac{∆E}{t}"

"P=\frac{\frac{1}{2}mv^2}{t}=\frac{mv^2}{2t}"

"t=\frac{mv^2}{2P}"